U=u(x1,x2)=(x1+2)^2(x2+3)3

find the marginal utility of each commodity


1
Expert's answer
2021-12-21T16:56:43-0500

marginal utilities:


"MU_1=\frac{\partial u(x_1,x_2)}{\partial x_1}=2(x_1+2)(x_2+3)^3"


"MU_1=\frac{\partial u(x_1,x_2)}{\partial x_2}=3(x_1+2)^2(x_2+3)^2"